1. Much of the demand for U.S. agricultural output has come from other countries. In 1998, the total demand (which is the sum of total domestic demand and total foreign demand) for wheat was Q_D=4200-250P. Of this, total domestic demand was Q_dom=2700-170P, and domestic supply was Q_S=1200+150P.

a. What is the demand function of the demand for wheat in foreign market Q_f?

b. What is the equilibrium price and quantity of the wheat market?

c. Suppose that China announces to decrease import of US agricultural products in response to the new tariffs imposed by the US government. The news prompts foreign demand for U.S. wheat to fall by 20 percent. What is the change in equilibrium price once the market clears again? What about the change in equilibrium quantity?

Now that Q_f shrinks by 20 percent, we have the new demand functions:

d. If the US government wants to restore the price of wheat to its original level, how much wheat should the government buy? How much would it cost?
Answer #1

Comin/ge Date O The demand furetien fol cwhal n yekat derived taking differerce fureign G0=400-250P -0 Aidem=2700-170p %3D 0-Comlin rs Sinve Imporitin Kulyls. cwbeal led to dalhe ninyponts Us demand wheno w o पि alcalamd IANuantit This Con be with byNOw pul Ne pa7.03/25 in f Dau Sp=3,900-234P =) 3,900 -234X7:03125 -)2,254.7 Ther new Eouilibium. Quiantib 8 22547 85 =3.900-2